# Bulk viscosity of nuclear matter with pions in the neutrino-trapped regime

### v1

@@ III.2 Weak interaction rates
$n_{B}\dfrac{\mathop{x_{\pi}}}{\mathop{dt}}$ $=-\lambda_{4}\delta\mu_{1}-\lambda_{5}\delta\mu_{2}+(\lambda_{3}+\lambda_{4}+\lambda_{5})\delta\mu_{3}.$ (18c)
@@ III.3 Susceptibilities
$\delta\mu_{i}$ $=A_{i}\dfrac{\delta n_{B}}{n_{B}}\cos{(\omega t)}+n_{B}\big{[}B_{i}\Re{(\delta x_{p}e^{i\omega t})}$
$+C_{i}\Re{(\delta x_{\mu}e^{i\omega t})}+D_{i}\Re{(\delta x_{\pi}e^{i\omega t})}\big{]}$
$=A_{i}\dfrac{\delta n_{B}}{n_{B}}\cos{(\omega t)}$ (22)
$+n_{B}\left[B_{i}\Re{(\delta x_{p})}+C_{i}\Re{(\delta x_{\mu})}+D_{i}\Re{(\delta x_{\pi})}\right]\cos{\left(\omega t\right)}$
$-n_{B}\left[B_{i}\Im{(\delta x_{p})}+C_{i}\Im{(\delta x_{\mu})}+D_{i}\Im{(\delta x_{\pi})}\right]\sin{\left(\omega t\right)},$
@@
$P$ $=P_{0}+\dfrac{\delta n_{B}}{n_{B}}\kappa_{T}^{-1}\cos{\left(\omega t\right)}+n_{B}\cos{\left(\omega t\right)}\times$ (23)
$\left[-A_{1}\Re{\left(\delta x_{p}\right)}+(A_{1}-A_{2})\Re{\left(\delta x_{\mu}\right)}+(A_{1}-A_{3})\Re{\left(\delta x_{\pi}\right)}\right]$
$+n_{B}\sin{\left(\omega t\right)}\times$
$\left[A_{1}\Im{\left(\delta x_{p}\right)}-(A_{1}-A_{2})\Im{\left(\delta x_{\mu}\right)}-(A_{1}-A_{3})\Im{\left(\delta x_{\pi}\right)}\right],$
@@
The susceptibilities $A_{i}$ describe how far the system is pushed out of beta equilibrium when the system is compressed if the chemical fractions are frozen. If $A_{i}=0$, then the system does not depart from beta equilibrium when the density is increased, and so with respect to $\delta\mu_{i}$, the system is conformal. However, the system may have other degrees of freedom that do change when the density is increased, and so we describe the situation where, say, $A_{1}=0$ as partially conformal. The Maxwell relations (Eqs. 21a-21c) indicate that the $A_{i}$ are related to the degree to which the pressure depends on a particular particle fraction. As will be seen later, generally bulk viscosity grows with the increased magnitude of the $A_{i}$. In particular, if all $A_{i}=0$, then bulk viscosity (of the type described here, related to beta equilibration) should be zero.
Figure 3: Susceptilities of the $A$ type from the EoS described in Sec. II plotted in merger conditions (top panels) and in supernovae conditions (bottom panels). The left column plots are at a density of $1n_{0}$, while the right column panels are at $1.6n_{0}$. The dashed lines correspond to the EoS without pions, while the solid lines indicate the EoS with pions. $A_{3}$ has no counterpart in a system without pions.
@@
Figure 4: Susceptilities of the $B$, $C$, and $D$ types from the EoS described in Sec. II plotted in merger conditions (top panels) and in supernovae conditions (bottom panels). The dashed lines correspond to an EoS without pions, while the solid lines indicate an EoS with pions. $D_{3}$ has no counterpart in a system without pions.
The susceptibilities $B$, $C$, and $D$ represent, at a fixed density and temperature, how far the system is pushed from chemical equilibrium when just one of the particle fractions is adjusted, keeping the others fixed. For example, $B_{1}$, $C_{1}$, and $D_{1}$ (roughly) represent the correlation of $\delta\mu_{1}$ with $x_{p}$, $x_{\mu}$, and $x_{\pi}$ respectively. The $B$, $C$, and $D$ susceptibilities convert $\lambda_{i}$ into $\gamma_{i}$, as we will see in a later section. We plot the susceptibilities $B_{i}$, $C_{i}$, and $D_{i}$ in Fig. 4. These susceptibilities, at least in the conditions encountered in this work, never cross zero. As we will see, while they do affect the strength of the maximum bulk viscosity, their most important role is their impact on the equilibration rates $\gamma_{i}$ which determine the location of the resonant maxima of the bulk viscosity.
@@ III.4 Bulk viscosity at finite $\lambda_{3}$
Combining Eqs. 9, 12, and 21a-21c, we find the bulk viscosity is given by
@@
To obtain $\Im{(\delta x_{i})}$, we plug Eq. 10 and III.3 into Eqs. 18a, 18b, and 18c and match the sine terms and the cosine terms. This yields six equations with six variables ($\Re{(\delta x_{i})}$ and $\Im{(\delta x_{i})}$ for $i=p,\mu,\pi$). The three real-part variables can be eliminated, yielding a set of 3 equations to be solved to obtain $\Im{(\delta x_{p})},\Im{(\delta x_{\mu})},\Im{(\delta x_{\pi})}$. Solving the system results in the expression for the bulk viscosity
@@ III.6 Bulk viscosity in the limit $\lambda_{3}\rightarrow\infty$
The limit where $n\leftrightarrow p+\pi^{-}$ occurs infinitely quickly is subtle. One might assume that $\delta\mu_{3}$ could simply be set to zero in Eq. 18a, 18b, and 18c. However, it is the product $\lambda_{3}\delta\mu_{3}$ that appears on the RHS of those equations, and even for very tiny deviations of $\delta\mu_{3}$ from zero, the product $\lambda_{3}\delta\mu_{3}$ could still be sizeable since $\lambda_{3}$ is very large. One must keep $\lambda_{3}$ finite throughout the calculation (up until Eq. 26) and then take the limit $\lambda_{3}\rightarrow\infty$. This was pointed out by [62, 63], and was applied to the case of hyperon bulk viscosity in [10].
@@ III.7 Partial bulk viscosities
Throughout this paper, we have called both $\lambda$ and $\gamma$ ”equilibration rates”. Each is useful in its own way. The difference in the rates, schematically, can be written $\overrightarrow{\Gamma}-\overleftarrow{\Gamma}\sim\lambda\left(\delta\mu-\delta\mu^{0}\right)=\lambda\delta\mu$ or $\overrightarrow{\Gamma}-\overleftarrow{\Gamma}\sim\gamma\left(n-n^{0}\right)$, where the superscript 0 denotes the beta equilibrium value. So, $\lambda$ describes the rate at which the chemical potential is forced to its beta equilibrium value, while $\gamma$ describes how quickly the particle fraction is forced to its beta equilibrium value by Le Chatelier’s principle. The susceptibilities of type $B$, $C$, and $D$ facilitate the conversion between $\gamma$ and $\lambda$. Note that both $\lambda_{i}$ and $\gamma_{i}$ are positive. $\gamma$ has a clearer physical meaning, as it has dimension one, and thus can be written as an inverse timescale and directly compared with the density oscillation frequency $\omega$.
@@
It is not clear to us that the prefactors in the $\lambda_{3}\rightarrow\infty$ limit can be written in terms of differences of compressibilities. But, it is useful to compare these expressions to their counterparts without pions (Eqs. 30a-30c) and with pions but without $\lambda_{3}\rightarrow\infty$ (Eqs. 59a-59f).
Figure 5: Partial bulk viscosities in matter with with $Y_{Le}=0.05$ and $Y_{L\mu}=0$ at $1n_{0}$ (left panel) and $1.6n_{0}$ (right panel). Results for pionless matter are in thin dotted lines. Results for matter with thermal pions with $\lambda_{3}=0$ (the conventional definition of partial bulk viscosities) are in thick dashed lines, and results with the “renormalized” partial bulk viscosities ($\lambda_{3}\rightarrow\infty$) are shown in thick solid lines. The total bulk viscosity - not the sum of the partial bulk viscosities (and discussed in Sec. IV) - is shown in grey.
In Fig. 5 we plot, just as an illustration, the partial bulk viscosities in the case of the pionless EoS ($\zeta_{1}^{\text{no }\pi}$, $\zeta_{2}^{\text{no }\pi}$, and $\zeta_{6}^{\text{no }\pi}$), for the EoS containing thermal pions where $\lambda_{3}$ is not sent to infinity (that is, the conventional definition of partial bulk viscosities in Appendix H), and for the EoS containing thermal pions but in the $\lambda_{3}\rightarrow\infty$ limit $\zeta_{i}^{\lambda_{3}\rightarrow\infty}$ ($i=1,2,4,5,6$). The full result for the bulk viscosity is shown in grey - dashed for the EoS without pions and solid for the EoS with pions. We see that most of the partial bulk viscosities display the predicted resonant structure, often with a conformal point at a temperature above that of the resonant maximum. The quantities $\zeta_{4}$ and $\zeta_{5}$ increase with temperature, because of the strong temperature dependence of the relevant susceptibilities (this overrides the traditional resonance structure, which assumes that the susceptibilities do not depend strongly on temperature [28]). We will refer back to this figure in the next section.
@@ IV.1 Equilibration rates and bulk viscosity
We consider neutron star merger conditions ($Y_{Le}=0.05,Y_{L\mu}=0$) first. The top left panel of Fig. 6 indicates that in the EoS without pions, all three rates $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{6}$ rise above $f=1\text{ kHz}$ in the temperature range of a few MeV. Indeed, one sees that the three partial bulk viscosities (Fig. 5) have resonant peaks in that temperature range. The full calculation for the bulk viscosity in pionless matter (Fig. 7) tracks the partial bulk viscosity $\zeta_{1}^{\text{no }\pi}$ for low temperature and through its resonant maximum at $T\approx 2$ MeV. As temperature increases, both $\gamma_{2}$ and $\gamma_{6}$ cross resonance at very similar temperatures, leading to a second resonant peak in the total bulk viscosity, that doesn’t quite overlap with any of the partial bulk viscosities. There is no third resonant peak, because the pionless EoS only has two equilibrating quantities, $x_{p}$ and $x_{\mu}$, so the third, slowest, rate is redundant. As temperature increases further, the bulk viscosity continues to decrease, since all particle fractions equilibrate very quickly compared to the millisecond timescale density oscillations. At a temperature of about 20 MeV, the system becomes partly conformal, because $A_{1}$ and $A_{2}$ pass through zero, and the bulk viscosity dips down to a sharp minimum, but not to zero (unlike the partial bulk viscosities, which do dip to zero). So, without pions, the neutrino-trapped $npe\mu$ bulk viscosity has two resonant peaks as a function of temperature, both at well below $T=10\text{ MeV}$, and then a pseudo-conformal dip at a temperature around 20 MeV.
@@
When pions are added to the EoS and their reactions are considered, the existing rates $\gamma_{1}$, $\gamma_{2}$, $\gamma_{6}$ change very little [^6] As can be seen in the top left panel of Fig. 6, including pions does change the rates at high temperature, where the pion population becomes substantial. But at these high temperatures, the bulk viscosity is very far off of resonance because the rates are much faster than the 1 kHz density oscillation, and so the bulk viscosity is small.. However, three new equilibration rates involving pions, $\gamma_{3}$, $\gamma_{4}$, and $\gamma_{5}$ enter the picture. The direct Urca (electron) rate $\gamma_{1}$ is still the fastest, and leads to the first resonant peak in the bulk viscosity, just as in the pionless case (though, in reality, this is the second resonant peak, because $\gamma_{3}$ would reach resonance at some temperature below the 1 MeV minimum studied here). The next resonant peak comes from $\gamma_{5}$, the pion decay to muons, which supersedes the reactions that dominate in the absence of pions. In fact, the resonance shifts to slightly lower temperatures. At temperatures just above that resonant peak ($T\gtrsim 4$ MeV), all three independent particle fractions are quick to equilibrate and thus the bulk viscosity decreases with temperature. While Fig. 6 shows other processes passing above the 1 kHz line (and thus Fig. 5 shows other resonant peaks at higher temperature, for example, $\zeta_{4}$), those never manifest in the full bulk viscosity calculation because they are redundant. The fastest three rates control the bulk viscosity in neutrino-trapped $npe\mu\pi^{-}$ matter.
@@ IV.2 Particle dynamics during an oscillation
Figure 9: In matter with $Y_{Le}=0.05,Y_{L\mu}=0$, the evolution of the three $\delta\mu$ quantities over the course of two oscillation periods (with oscillation amplitude $\delta n_{B}=0.03$). Dashed lines denote matter without pions, solid lines denote matter including thermal pions. Rows 1, 2, and 3 are at temperatures of 2, 5, and 10 MeV, respectively. The left column is in matter at $1n_{0}$ and the right column is for matter at $1.6n_{0}$. Note that the y axes do not have the same scale. The density oscillation $n_{B}(t)$ is overlaid in pink so that one can tell the degree to which the particle fractions are in phase with the density oscillation.
@@
Figure 10: In matter with $Y_{Le}=0.05,Y_{L\mu}=0$, the evolution of the three independent particle fractions $x_{p}$, $x_{\mu}$, and $x_{\pi}$ (we plot the fractional deviation from the beta equilibrium value of each particle fraction) over the course of two oscillation periods (with oscillation amplitude $\delta n_{B}=0.03$). Dashed lines denote matter without pions, solid lines denote matter including thermal pions. Rows 1, 2, and 3 are at temperatures of 2, 5, and 10 MeV, respectively. The left column depicts matter at $1n_{0}$ and the right column shows matter at $1.6n_{0}$. The density oscillation $n_{B}(t)$ is overlaid in pink so that one can tell the degree to which the particle fractions are in phase with the density oscillation.
@@
In Fig. 9 we plot the evolution of $\delta\mu_{1}$, $\delta\mu_{2}$, and $\delta\mu_{3}$ over the course of the harmonic density oscillation. When these quantities depart from zero, the matter is no longer in chemical equilibrium. In the limit $\lambda_{3}\rightarrow\infty$, $\delta\mu_{3}$ is forced to zero, so the reaction $n\leftrightarrow p+\pi^{-}$ is always in equilibrium.
@@
At $1n_{0}$ and $T=2\text{ MeV}$ (top left panel of Fig. 9), both $\delta\mu_{1}(t)$ and $\delta\mu_{2}(t)$ oscillate up to a couple of MeV out of beta equilibrium. The addition of pions into the system at this low temperature and density makes no difference because the pion number is very low. At its maximum, the quantity $\delta\mu_{2}$ exceeds 2 MeV, and is no longer small compared to the temperature, which indicates that the subthermal approximation taken here loses validity - suprathermal corrections [28, 65] should be taken into account at these low temperatures. As the temperature rises to 5 MeV, the chemical reaction rates grow quickly and now the direct Urca process (with an electron) is fast enough to keep $\delta\mu_{1}$ very close to chemical equilibrium throughout the oscillation. The quantity $\delta\mu_{2}$ is still pushed by about 1 MeV out of equilibrium, as the muon population is too low for the direct Urca processes with a muon ($\gamma_{2}$) (or muon-electron conversion $\gamma_{6}$) to equilibrate it efficiently. When pions are included in the EoS, their population is sufficiently high at $T=5\text{ MeV}$ that $\delta\mu_{2}$ is kepts closer to equilibrium than without the pions present. Finally, at $T=10\text{ MeV}$, the system (with or without pions) remains extremely close to beta equilibrium (note the y-axis scale in the bottom left panel of Fig. 9) throughout the density oscillation, indicating that the bulk viscosity will be small at this high temperature.

### v3

@@ III.2 Weak interaction rates
$n_{B}\dfrac{\mathop{dx_{\pi}}}{\mathop{dt}}$ $=-\lambda_{4}\delta\mu_{1}-\lambda_{5}\delta\mu_{2}+(\lambda_{3}+\lambda_{4}+\lambda_{5})\delta\mu_{3}.$ (20c)
@@ III.3 Susceptibilities
$\delta\mu_{i}$ $=A_{i}\dfrac{\delta n_{B}}{n_{B}}\cos{(\omega t)}+n_{B}\left[B_{i}\Re{(\delta x_{p}e^{i\omega t})}+C_{i}\Re{(\delta x_{\mu}e^{i\omega t})}+D_{i}\Re{(\delta x_{\pi}e^{i\omega t})}\right]$ (24)
$=A_{i}\dfrac{\delta n_{B}}{n_{B}}\cos{(\omega t)}+n_{B}\left[B_{i}\Re{(\delta x_{p})}+C_{i}\Re{(\delta x_{\mu})}+D_{i}\Re{(\delta x_{\pi})}\right]\cos{\left(\omega t\right)}-n_{B}\left[B_{i}\Im{(\delta x_{p})}+C_{i}\Im{(\delta x_{\mu})}+D_{i}\Im{(\delta x_{\pi})}\right]\sin{\left(\omega t\right)},$
@@
$P$ $=P_{0}+\dfrac{\delta n_{B}}{n_{B}}\kappa_{T}^{-1}\cos{\left(\omega t\right)}+n_{B}\cos{\left(\omega t\right)}\left[-A_{1}\Re{\left(\delta x_{p}\right)}+(A_{1}-A_{2})\Re{\left(\delta x_{\mu}\right)}+(A_{1}-A_{3})\Re{\left(\delta x_{\pi}\right)}\right]$
$+n_{B}\sin{\left(\omega t\right)}\left[A_{1}\Im{\left(\delta x_{p}\right)}-(A_{1}-A_{2})\Im{\left(\delta x_{\mu}\right)}-(A_{1}-A_{3})\Im{\left(\delta x_{\pi}\right)}\right],$ (25)
@@
The susceptibilities $A_{i}$ describe how far the system is pushed out of beta equilibrium when the system is compressed if the chemical fractions are frozen. If $A_{i}=0$, then the system does not depart from beta equilibrium when the density is increased, and so with respect to $\delta\mu_{i}$, the system is conformal. However, the system may have other degrees of freedom that do change when the density is increased, and so we describe the situation where, say, $A_{1}=0$ as partially conformal. The Maxwell relations (Eqs. 23a-23c) indicate that the $A_{i}$ are related to the degree to which the pressure depends on a particular particle fraction. As will be seen later, generally bulk viscosity grows with the increased magnitude of the $A_{i}$. In particular, if all $A_{i}=0$, then bulk viscosity (of the type described here, related to chemical equilibration) should be zero.
Figure 3: Susceptibilities of the $A$ type from the EoS described in Sec. II plotted in merger conditions (top panels) and in supernovae conditions (bottom panels). The left column plots are at a density of $1n_{0}$, while the right column panels are at $1.6n_{0}$. The dashed lines correspond to the EoS without pions, while the solid lines indicate the EoS with pions. $A_{3}$ has no counterpart in a system without pions.
@@
Figure 4: Susceptibilities of the $B$, $C$, and $D$ types from the EoS described in Sec. II plotted in merger conditions (top panels) and in supernovae conditions (bottom panels). The dashed lines correspond to an EoS without pions, while the solid lines indicate an EoS with pions. $D_{3}$ has no counterpart in a system without pions.
The susceptibilities $B$, $C$, and $D$ represent, at a fixed density and temperature, how far the system is pushed from chemical equilibrium when just one of the particle fractions is adjusted, keeping the others fixed. For example, $B_{1}$, $C_{1}$, and $D_{1}$ (roughly) represent the degree of dependence of $\delta\mu_{1}$ on $x_{p}$, $x_{\mu}$, and $x_{\pi}$ respectively. The $B$, $C$, and $D$ susceptibilities convert $\lambda_{i}$ into $\gamma_{i}$, as we will see in a later section. We plot the susceptibilities $B_{i}$, $C_{i}$, and $D_{i}$ in Fig. 4. These susceptibilities, at least in the conditions encountered in this work, never cross zero. As we will see, while they do affect the strength of the maximum bulk viscosity, their most important role is their impact on the equilibration rates $\gamma_{i}$ which determine the location of the resonant maxima of the bulk viscosity.
@@ III.4 Bulk viscosity at finite $\lambda_{3}$
Combining Eqs. 11, 14, and 23a-23c, we find the bulk viscosity is given by
@@
To obtain $\Im{(\delta x_{i})}$, we plug Eq. 12 and 24 into Eqs. 20a, 20b, and 20c and match the sine terms and the cosine terms. This yields six equations with six variables ($\Re{(\delta x_{i})}$ and $\Im{(\delta x_{i})}$ for $i=p,\mu,\pi$). The three real-part variables can be eliminated, yielding a set of 3 equations to be solved to obtain $\Im{(\delta x_{p})},\Im{(\delta x_{\mu})},\Im{(\delta x_{\pi})}$. Solving the system results in the expression for the bulk viscosity
@@ III.6 Bulk viscosity in the limit $\lambda_{3}\rightarrow\infty$
The limit where $n\leftrightarrow p+\pi^{-}$ occurs infinitely quickly is subtle. One might assume that $\delta\mu_{3}$ could simply be set to zero in Eq. 20a, 20b, and 20c. However, it is the product $\lambda_{3}\delta\mu_{3}$ that appears on the RHS of those equations, and even for very tiny deviations of $\delta\mu_{3}$ from zero, the product $\lambda_{3}\delta\mu_{3}$ could still be sizeable since $\lambda_{3}$ is very large. One must keep $\lambda_{3}$ finite throughout the calculation (up until Eq. 28) and then take the limit $\lambda_{3}\rightarrow\infty$. This was pointed out by [63, 64], and was applied to the case of hyperon bulk viscosity in [10].
@@ III.7 Partial bulk viscosities
Throughout this paper, we have called both $\lambda$ and $\gamma$ ”equilibration rates”. Each is useful in its own way. The difference in the rates, schematically, can be written $\overrightarrow{\Gamma}-\overleftarrow{\Gamma}\sim\lambda\left(\delta\mu-\delta\mu^{0}\right)=\lambda\delta\mu$ or $\overrightarrow{\Gamma}-\overleftarrow{\Gamma}\sim\gamma\left(n-n^{0}\right)$, where the superscript 0 denotes the beta equilibrium value. So, $\lambda$ describes the rate at which the chemical potential difference is forced to its beta equilibrium value (zero), while $\gamma$ describes how quickly the particle fraction is forced to its beta equilibrium value by Le Chatelier’s principle. The susceptibilities of type $B$, $C$, and $D$ facilitate the conversion between $\gamma$ and $\lambda$. Note that both $\lambda_{i}$ and $\gamma_{i}$ are positive. $\gamma$ has a clearer physical meaning, as it has dimension one, and thus can be written as an inverse timescale and directly compared with the density oscillation frequency $\omega$.
@@
It is not clear to us that the prefactors in the $\lambda_{3}\rightarrow\infty$ limit can be written in terms of differences of compressibilities. But, it is useful to compare these expressions to their counterparts without pions (Eqs. 32a-32c) and with pions but without $\lambda_{3}\rightarrow\infty$ (Eqs. 61a-61f).
Figure 5: Partial bulk viscosities in matter with $Y_{Le}=0.05$ and $Y_{L\mu}=0$ at $1n_{0}$ (left column) and $1.6n_{0}$ (right column). Results for pionless matter are in thin dotted lines. The top row depicts the conventional definition of partial bulk viscosities, where $\lambda_{3}$ is set to zero. The bottom row depicts the “renormalized” definition of partial bulk viscosities where $\lambda_{3}$ is sent to infinity. The total bulk viscosity - not the sum of the partial bulk viscosities (and discussed in Sec. IV) - is shown with a thick grey line. For clarity, for this figure the x-axes are displayed with a log scale.
In Fig. 5 we plot, just as an illustration, the partial bulk viscosities in the case of the pionless EoS ($\zeta_{1}^{\text{no }\pi}$, $\zeta_{2}^{\text{no }\pi}$, and $\zeta_{6}^{\text{no }\pi}$), for the EoS containing thermal pions where $\lambda_{3}$ is not sent to infinity (that is, the conventional definition of partial bulk viscosities in Appendix H), and for the EoS containing thermal pions but in the $\lambda_{3}\rightarrow\infty$ limit $\zeta_{i}^{\lambda_{3}\rightarrow\infty}$ ($i=1,2,4,5,6$). The full result for the bulk viscosity is shown with a thick grey line - dashed for the EoS without pions and solid for the EoS with pions. We see that most of the partial bulk viscosities display the predicted resonant structure, often with a conformal point at a temperature above that of the resonant maximum. The quantities $\zeta_{4}$ and $\zeta_{5}$ increase with temperature, because of the strong temperature dependence of the relevant susceptibilities (this overrides the traditional resonance structure, which assumes that the susceptibilities do not depend strongly on temperature [28]). The total bulk viscosity (with pions) seems to track the “renormalized” ($\lambda_{3}\rightarrow\infty$) partial bulk viscosities much better than the traditional ($\lambda_{3}=0$) versions. We will refer back to this figure in the next section.
@@ IV.1 Equilibration rates and bulk viscosity
We consider neutron star merger conditions ($Y_{Le}=0.05,Y_{L\mu}=0$) first. The top left panel of Fig. 6 indicates that in the EoS without pions, all three rates $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{6}$ rise above $f=1\text{ kHz}$ in the temperature range of a few MeV. Indeed, one sees that the three partial bulk viscosities (Fig. 5) have resonant peaks in that temperature range. The full calculation for the bulk viscosity in pionless matter (Fig. 7) tracks the partial bulk viscosity $\zeta_{1}^{\text{no }\pi}$ for low temperature and through its resonant maximum at $T\approx 2$ MeV. As temperature increases, both $\gamma_{2}$ and $\gamma_{6}$ cross resonance at very similar temperatures, leading to a second resonant peak in the total bulk viscosity, that doesn’t quite overlap with any of the partial bulk viscosities. There is no third resonant peak, because the pionless EoS only has two equilibrating quantities, $x_{p}$ and $x_{\mu}$, so the third, slowest, rate is redundant. As temperature increases further, the bulk viscosity continues to decrease, since all particle fractions equilibrate very quickly compared to the millisecond timescale density oscillations. At a temperature of about 20 MeV, the system becomes partly conformal, because $A_{1}$ and $A_{2}$ pass through zero, and the bulk viscosity dips down to a sharp minimum, but not to zero (unlike the partial bulk viscosities, which do dip to zero). So, without pions, the neutrino-trapped $npe\mu$ bulk viscosity has two resonant peaks as a function of temperature, both at well below $T=10\text{ MeV}$, and then a partial conformal dip at a temperature around 20 MeV.
@@
When pions are added to the EoS and their reactions are considered, the existing rates $\gamma_{1}$, $\gamma_{2}$, $\gamma_{6}$ change very little [^7] As can be seen in the top left panel of Fig. 6, including pions does change the rates at high temperature, where the pion population becomes substantial. But at these high temperatures, the bulk viscosity is very far off of resonance because the rates are much faster than the 1 kHz density oscillation, and so the bulk viscosity is small.. However, three new equilibration rates involving pions, $\gamma_{3}$, $\gamma_{4}$, and $\gamma_{5}$ enter the picture. The direct Urca (electron) rate $\gamma_{1}$ is still the fastest (besides $\gamma_{3}$), and leads to the first resonant peak in the bulk viscosity, just as in the pionless case (though, in reality, this is the second resonant peak, because $\gamma_{3}$ would reach resonance at some temperature below the 1 MeV minimum studied here). The next resonant peak comes from $\gamma_{5}$, the pion decay to muons, which supersedes the reactions that dominate in the absence of pions. In fact, the resonance shifts to slightly lower temperatures. At temperatures just above that resonant peak ($T\gtrsim 4$ MeV), all three independent particle fractions are quick to equilibrate and thus the bulk viscosity decreases with temperature. While Fig. 6 shows other processes passing above the 1 kHz line (and thus Fig. 5 shows other resonant peaks at higher temperature, for example, $\zeta_{4}$), those never manifest in the full bulk viscosity calculation because they are redundant. The fastest three rates control the bulk viscosity in neutrino-trapped $npe\mu\pi^{-}$ matter.
@@ IV.2 Particle dynamics during an oscillation
Figure 9: In matter with $Y_{Le}=0.05,Y_{L\mu}=0$, the evolution of the three $\delta\mu$ quantities over the course of two oscillation periods (with oscillation amplitude $\delta n_{B}=0.03n_{0}$). Dashed lines denote matter without pions, solid lines denote matter including thermal pions. Rows 1, 2, and 3 are at temperatures of 2, 5, and 10 MeV, respectively. The left column is in matter at $1n_{0}$ and the right column is for matter at $1.6n_{0}$. Note that the y axes do not have the same scale. The density oscillation $n_{B}(t)$ is overlaid in pink so that one can tell the degree to which the $\delta\mu_{i}$ are in phase with the density oscillation.
@@
Figure 10: In matter with $Y_{Le}=0.05,Y_{L\mu}=0$, the evolution of the three independent particle fractions $x_{p}$, $x_{\mu}$, and $x_{\pi}$ over the course of two oscillation periods (with oscillation amplitude $\delta n_{B}=0.03n_{0}$). Dashed lines denote matter without pions, solid lines denote matter including thermal pions. Rows 1, 2, and 3 are at temperatures of 2, 5, and 10 MeV, respectively. The left column depicts matter at $1n_{0}$ and the right column shows matter at $1.6n_{0}$. The density oscillation $n_{B}(t)$ is overlaid in pink so that one can tell the degree to which the particle fractions are in phase with the density oscillation.
@@
In Fig. 9 we plot the evolution of $\delta\mu_{1}$, $\delta\mu_{2}$, and $\delta\mu_{3}$ over the course of the harmonic density oscillation. These quantities are obtained from Eq. 24, where the real and imaginary parts of $\delta x_{i}$ are obtained from the solution of the system of equations described after Eq. 27. When any of the $\delta\mu_{i}$ departs from zero, the matter is no longer in chemical equilibrium. In the limit $\lambda_{3}\rightarrow\infty$, $\delta\mu_{3}$ is forced to zero, so the reaction $n\leftrightarrow p+\pi^{-}$ is always infinitesimally close to equilibrium.
@@
At $1n_{0}$ and $T=2\text{ MeV}$ (top left panel of Fig. 9), both $\delta\mu_{1}(t)$ and $\delta\mu_{2}(t)$ oscillate up to a couple of MeV out of beta equilibrium. The addition of pions into the system at this low temperature and density makes no difference because the pion number is very low. At its maximum, the quantity $\delta\mu_{2}$ exceeds 2 MeV, and is no longer small compared to the temperature, which indicates that the subthermal approximation taken here loses validity - suprathermal corrections [28, 66] should be taken into account at these low temperatures. As the temperature rises to 5 MeV, the chemical reaction rates grow quickly and now the direct Urca process (with an electron) is fast enough to keep $\delta\mu_{1}$ very close to chemical equilibrium throughout the oscillation. The quantity $\delta\mu_{2}$ is still pushed by about 1 MeV out of equilibrium, as the muon population is too low for the direct Urca processes with a muon ($\gamma_{2}$) (or muon-electron conversion $\gamma_{6}$) to equilibrate it efficiently. When pions are included in the EoS, their population is sufficiently high at $T=5\text{ MeV}$ that $\delta\mu_{2}$ is kept closer to equilibrium than without the pions present. Finally, at $T=10\text{ MeV}$, the system (with or without pions) remains extremely close to beta equilibrium (note the y-axis scale in the bottom left panel of Fig. 9) throughout the density oscillation, indicating that the bulk viscosity will be small at this high temperature.